OpenStudy (anonymous):

when i have to multiply for example 4y^2 .........y being -4 it give me -16^2 but ^2 means 16 x16 ?

OpenStudy (anonymous):

so that means when I multiply it will b multiplied with the exponent

OpenStudy (campbell_st):

not quite squaring a negative gives a positive so you are looking at \[4 \times (-4)^2\] remember order of operation.....

OpenStudy (anonymous):

pemdas right?

OpenStudy (campbell_st):

the way the question is written, the exponent of 2 is only operating on y

OpenStudy (anonymous):

y^2= y*y -4*-4 However be very careful with calculators sometimes they do give a -